The Tropical Breezes Company expects their dividends and earnings to grow at a constant rate of 4% a year into the foreseeablA)

A)

B)

C)

pls type not write, and calculations need to be shown w/ steps to TI BAII PLUS CALCULATOR
Answer #1

Recently paid dividend , DO = $2.75 Growth rate, g = 4% Rate of return, RR = 10% D1 = DO * (1+g) = $2.75*(1 + 4%) = $2.86 a)